Working together Lolita and Doug can paint a shed in 2 hours. If Doug can paint the shed in 8 hours, how long will it take Lolita to paint the shed by herself?
Answer by richwmiller(17219)

2/L+2/8=1
2/L=3/4
8=3L
8/3=L